OpenStudy (anonymous):

I'll write testimony, fan and medal for anyone would could help be solve this step by step!! Arthur has decided to start saving for a new computer. His money is currently in a piggy bank at home, modeled by the function s(x) = 85. He was told that he could do the laundry for the house and his allowance would be a(x) = 10(x − 1), where x is measured in weeks. Explain to Arthur how he can create a function that combines the two, and describe any simplification that can be done.

OpenStudy (perl):

you can add the two functions to get the total money s(x) + a(x) = 85 + 10(x-1)

OpenStudy (perl):

now simplified s(x) + a(x) is = 85 + 10x - 10 = 75 + 10 x

OpenStudy (perl):

distribute

OpenStudy (anonymous):

@perl could you help me explain it, I know what you're talking about, which is awesome for me considering I suck, I just don't have words to describe. and @Shawn1324 giving me the answer did nothing.... this isn't multiple choice. I have to explain myself.

OpenStudy (perl):

you have two sources of income. the piggy bank s(x) = 85 , and the allowance a(x)= 10(x-1) Notice that the piggy bank is a constant function, since the money is just sitting in the piggy bank.

OpenStudy (anonymous):

am I suppose to divide 10 to both sides? 10x = 75? @perl